In a hall, there are 180 seats. The number of seats in each row is 8 more than the number of rows. Find the number of rows
Answer by mananth(16946)

rows be x
seats in each row = x+8
seats = 180
x(x+8)=180
x^2+8x=180
x^2+8x-180=0
x^2+18x-10x-180=0
x(x+18)-10(x+18)=0
(x+18)(x-10)=0
x= 10 number of rows